Write the equation y - 2 = 2(x - 3) in standard form.
Answer by stanbon(75887)

Write the equation y - 2 = 2(x - 3) in standard form.
y - 2 = 2x-6
y = 2x-4
----
Standard form:
2x-y = 4
-------------------
Cheers,
Stan H.